# Enhanced nuclear Schiff and electric dipole moments in nuclei with an octupole deformation

### v1

@@ Abstract
In nuclei with octupole deformation, close opposite-parity rotational states with identical spin are mixed by $T$,$P$-violating nuclear forces. This mixing polarises the nuclear axis along the nuclear spin, allowing moments from the intrinsic frame to manifest in the laboratory frame, provided the nuclear spin $I$ is sufficiently large. This mechanism may be extended to nuclei with a soft octupole vibration mode.
Using half-life data for $E1$ transitions from the NuDat database, we calculate the intrinsic nuclear EDM $d_{\text{int}}$ for a range of nuclei theorised to exhibit octupole deformation or soft octupole vibration. From these values, we independently estimate the intrinsic nuclear Schiff moment $S_{\text{int}}$ and the octupole deformation parameter $\beta_{3}$. Finally, we compare the magnitude of these collective moments in the laboratory frame with the contributions from valence nucleons, providing an estimate of the nuclear EDM and Schiff moment components unrelated to octupole deformation.
@@ I Introduction
Refs. [13, 14, 23, 24] have performed numerical calculations of the Schiff moments and estimates of the atomic EDM produced by the electrostatic interaction between electrons and these moments for 221Fr, 223Fr, 223Ra, 223Rn, 225Ra, 225Ac and 229Pa. The short lifetime of these nuclei proves to be an issue when it comes to experiments. Such experiments for 223Rn and 225Ra have been considered by various experimental groups [25, 26, 27]. Despite the Schiff moment enhancement, the 225Ra EDM measurements presented in Refs. [26, 27] have not reached the sensitivity to the $T$,$P$-violating interaction as the Hg EDM experiment [28]. These experiments are hindered by the instability of 225Ra (which has a half-life of 15 days) and a relatively small number of atoms. Ref. [29] extended the list of candidate nuclei to include stable isotopes 153Eu, 155Gd, 161Dy and 163Dy, as well as long lifetime nuclei 153Sm,165Er,225Ac, 227Ac, 229Th, 229Pa, 233U, 235U, 237Np and 239Pu.
@@ II Nuclear Shape and Intrinsic Moments
$c_{V}=1-\left(\frac{1}{\sqrt{4\pi}}\right)\sum_{l=1}^{\infty}\beta_{l}^{2}\,,$ (4)
@@
$\frac{\rho_{p}(\vec{r})-\rho_{n}(\vec{r})}{\rho_{p}(\vec{r})+\rho_{n}(\vec{r})}=-\frac{1}{4C}V_{\text{Coul}}(\vec{r})\,,$ (7)
where $V_{\text{Coul}}(\vec{r})$ is the Coulomb potential created by the distribution of the proton density and $C$ is the volume symmetry-energy coefficient of the liquid drop model. To Lowest order, the proton density may be written as [43]
@@ III Calculation of Enhanced Nuclear Moments
where $K_{i}$ and $K_{f}$ correspond to the projection of the nuclear spin $I$ onto the rotation axis of the initial and final states respectively. As such, combining Eqs. (15) and (16) (considering electrically induced transitions with a multipolarity of $l_{\gamma}=1$) we yield the following expression for the intrinsic nuclear EDM
@@
This point may need clarification. Static octupole deformation is in fact a theoretical construction meaning that the minimum in energy is achieved in the case of an octupole deformation. The potential energy in this case is actually a double potential well with two equal minima (a bump on one side and a bump on the other side). In a quantum oscillator with a double potential well, the stationary states are linear combinations of the states with a particle in the first potential well and a particle in the second potential well. In this case $\expectationvalue{x}=0$, but $\expectationvalue{x^{2}}\neq 0$ and is relatively large, of the order of the distance between the potential energy minima. This is why there is no principal difference between “static” octupole deformation and soft octupole vibrations with $\expectationvalue{\beta_{3}}=0$ and a large $\expectationvalue{\beta_{3}^{2}}$.
@@
where $g$ is the strong $\pi$-meson - nucleon interaction constant and $\bar{g}_{0}$, $\bar{g}_{1}$, $\bar{g}_{2}$ are the $\pi$-meson - nucleon $CP$-violating interaction constants in the isotopic channels $T$ = 0, 1, 2. Given this, we may express the results for the nuclear EDM and Schiff moment in terms of more fundamental parameters such as the QCD $\theta$-term constant $\bar{\theta}$ using the relation $|g\bar{g}_{0}|$ = 0.21 $|\bar{\theta}|$ and $|g\bar{g}_{1}|$ = -0.046 $|\bar{\theta}|$ [47, 48, 49]. Using these results, we obtain
@@
The orders of magnitude of the Schiff moments are consistent with those presented in e.g. Refs. [29, 50]. Recent studies in Refs. [51, 52] have estimated the Schiff moment of 153Eu using different nuclear models that do not assume a static octupole deformation. The result calculated in Ref. [51] is two times smaller than our present result, while the calculations of Ref. [52] predict an even smaller value.
@@
Calculations were not be able to be performed for 163Dy, 223Rn, 229Th and 235U (which are also theorised to have an octupole deformed nucleus or a soft octupole vibration mode) due to inconclusive $E1$ transition data in Ref. [45]. For a similar reason, we take the value of $d_{\text{int}}$ for 223Fr and 229Pa from other experiments, see Table 1.
@@
87 221Fr 5/2 0.18 15 0.16 0.41 0.34
87 223Fr 111$\lvert d_{\text{int}}\rvert$ obtained from Refs. [53, 54]. 3/2 0.24 19 0.18 0.76 0.60
@@ V Effects of Enhanced nuclear moments in Molecules
In this work, we have performed a series of calculations characterising the enhancement of nuclear moments provided by the so called octupole mechanism. Octupole deformed nuclei have a low-lying energy level near the ground state, with opposing parity and identical spin. As such, the $T,P$-odd nuclear moments which are produced through the mixing of these doublet states by $T$,$P$-violating nuclear forces should be enhanced. This enhancement may also be extended to nuclei which are theorised to have a soft octupole vibration mode. Therefore, we have calculated the intrinsic nuclear electric dipole moment in a range of nuclei which are theorised to exhibit either a static octupole deformation or a soft octupole vibration mode, using half-life data for $E1$ transitions. These values were then used to present independent estimates of the octupole deformation parameter of each nucleus, as well as the intrinsic Schiff moment. The mixing of the rotational doublet states in these nuclei polarises the orientation of the nuclear axis along the nuclear spin, resulting in the manifestation of these intrinsic moments in the laboratory frame. This allows us to estimate the enhanced laboratory frame values of the nuclear EDM and Schiff moment in these nuclei. Our calculations indicate an enhancement of approximately 2-3 orders of magnitude of this collective Schiff moment when compared to the valence nucleon contribution to the nuclear Schiff moment in spherical nuclei. The nuclear EDM may be enhanced up to one order of magnitude.
@@
Our calculation of the nuclear EDM and Schiff moment has implications on the search for ultralight dark matter which produces an oscillating neutron EDM [30], nuclear EDM and Schiff moment [31]. In order to yield enhanced results, experiments measuring the axion field (or axion-like particle field) induced Schiff moment are seeking to probe crystals containing nuclei which benefit from the octupole mechanism. In particular, our calculation of the Schiff moment in 153Eu may be relevant to experimental groups currently probing crystals containing this isotope such as Refs. [38, 63].
@@ Acknowledgements
We are grateful to Alex Sushkov and Oleg Sushkov for useful discussions. The work was supported by the Australian Research Council Grants No. DP230101058 and No. DP200100150.

### v3

@@ Abstract
In nuclei with octupole deformation, close opposite-parity rotational states with identical spin are mixed by $T$,$P$-violating nuclear forces. This mixing polarises the nuclear axis along the nuclear spin, allowing moments from the intrinsic frame to manifest in the laboratory frame, provided the nuclear spin $I$ is sufficiently large.
Using half-life data for $E1$ transitions from the NuDat database, we calculate the intrinsic nuclear EDM $d_{\text{int}}$ for a range of nuclei theorised to exhibit octupole deformation. From these values, we independently estimate the intrinsic nuclear Schiff moment $S_{\text{int}}$ and the octupole deformation parameter $\beta_{3}$. Finally, we compare the magnitude of these collective moments in the laboratory frame with the contributions from valence nucleons, providing an estimate of the nuclear EDM and Schiff moment components unrelated to octupole deformation. The uncertainty of our estimates may exceed a factor of 10.
@@ I Introduction
Refs. [13, 14, 23, 24] have performed numerical calculations of the Schiff moments and estimates of the atomic EDM produced by the electrostatic interaction between electrons and these moments for 221Fr, 223Fr, 223Ra, 223Rn, 225Ra, 225Ac and 229Pa. The short lifetime of these nuclei proves to be an issue when it comes to experiments. Such experiments for 223Rn and 225Ra have been considered by various experimental groups [25, 26, 27]. Despite the Schiff moment enhancement, the 225Ra EDM measurements presented in Refs. [26, 27] have not reached the sensitivity to the $T$,$P$-violating interaction as the Hg EDM experiment [28]. These experiments are hindered by the instability of 225Ra (which has a half-life of 15 days) and a relatively small number of atoms. Refs. [29, 30] extended the list of candidate nuclei to include stable isotopes 153Eu, 155Gd, 161Dy and 163Dy, as well as long lifetime nuclei 153Sm,165Er,225Ac, 227Ac, 229Th, 229Pa, 233U, 235U, 237Np and 239Pu.
Using this list of candidates as a guide, we use the half-life data from the NuDat database to estimate the nuclear EDM in the hydrodynamic liquid-drop model, with a particle coupled to a hydrodynamic core. These values are then used to estimate the nuclear Schiff moment and the octupole deformation parameter. The calculations presented in this paper may serve as a guide for future studies of nuclei theorised to have octupole deformation and may stimulate the undertaking of sophisticated nuclear many-body calculations (as well as atomic, molecular and solid state calculations) for new candidate nuclei. Such many-body calculations are resource-intensive and have thus far been limited to 225Ra [23] (and potentially 153Eu [31]) over nearly three decades since the original proposal in [13].
@@ II Nuclear Shape and Intrinsic Moments
$c_{V}=1-\left(\frac{1}{4\pi}\right)\sum_{l=1}^{\infty}\beta_{l}^{2}\,,$ (4)
@@
$\frac{\rho_{p}(\vec{r})-\rho_{n}(\vec{r})}{\rho_{p}(\vec{r})+\rho_{n}(\vec{r})}=-\frac{1}{4C}eV_{\text{Coul}}(\vec{r})\,,$ (7)
where $V_{\text{Coul}}(\vec{r})$ is the Coulomb potential created by the distribution of the proton density and $C$ is the volume symmetry-energy coefficient of the liquid drop model. We note that this equality is a kind of static equilibrium condition. In the case of octupole vibrations, it may lead to a significant error, especially for oscillations with frequencies of 1 MeV or higher.
To lowest order, the proton density may be written as [45]
@@
We note that the expressions presented in this section are based on the hydrodynamic (liquid-drop) model and do not include shell corrections. Refs. [45, 47] show that the shell correction to the dipole moment (13) is of the same order of magnitude as $d_{\text{int}}$. The calculations presented in the following sections are only first estimates and may be used in the consideration of future theoretical and experimental studies of enhanced nuclear moments.
@@ III Calculation of Enhanced Nuclear Moments
where $K_{i}$ and $K_{f}$ correspond to the projection of the nuclear spin $I$ onto the rotation axis of the initial and final states respectively. Note that in the case of octupole vibrations, $d_{\text{int}}$ may be regarded as a transition moment (between $0^{-}$ and $1^{-}$ phonon states) rather than an expectation value.
As such, combining Eqs. (15) and (16) (considering electrically induced transitions with a multipolarity of $l_{\gamma}=1$) we yield the following expression for the intrinsic nuclear EDM
@@
This point may require clarification. Static octupole deformation is, in fact, a theoretical construct, referring to a situation where the minimum energy configuration involves an octupole deformation. The potential energy in this case forms a double-well structure with two equal minima (representing a bump on one side and a bump on the other). In a quantum oscillator with such a double potential well, the stationary states are linear combinations of the states where the particle resides in either well. As a result, $\expectationvalue{x}=0$, but $\expectationvalue{x^{2}}\neq 0$ and remains relatively large, on the order of the separation between the potential minima. This explains why there is no fundamental difference between ”static” octupole deformation and soft octupole vibrations, where $\expectationvalue{\beta_{3}}=0$ but $\expectationvalue{\beta_{3}^{2}}$ is large.
However, there is a significant difference in the quantitative approach. When the potential energy does not have a sufficiently deep minimum at a non-zero $\beta_{3}$, Eq. (16) could be applicable under the adiabatic approximation for octupole oscillations, where the rotational frequency is much larger than the octupole vibration frequency. In such cases, we can treat the rotation as occurring for a fixed deformation parameter and then replace the variable $\beta_{3}^{2}$ with its average value $\expectationvalue{\beta_{3}^{2}}$.
In reality however, the rotational frequency is smaller than the octupole vibration frequency. As a result, the use of Eq. (16) within our semi-empirical approach provides only an upper limit for the EDM and Schiff moments.
@@
where $g$ is the strong $\pi$-meson - nucleon interaction constant and $\bar{g}_{0}$, $\bar{g}_{1}$, $\bar{g}_{2}$ are the $\pi$-meson - nucleon $CP$-violating interaction constants in the isotopic channels $T$ = 0, 1, 2. Given this, we may express the results for the nuclear EDM and Schiff moment in terms of more fundamental parameters such as the QCD $\theta$-term constant $\bar{\theta}$ using the relation $g\bar{g}_{0}=0.21\ \bar{\theta}$ and $g\bar{g}_{1}=-0.046\ \bar{\theta}$ [50, 51, 52]. Using these results, we obtain
@@
The orders of magnitude of the Schiff moments are consistent with those presented in e.g. Refs. [30, 53]. Recent studies in Refs. [54, 31] have estimated the Schiff moment of 153Eu using different nuclear models that do not assume a static octupole deformation. The result calculated in Ref. [54] is two times smaller than our present result, while the calculations of Ref. [31] predict an even smaller value. Approximations in the evaluation of the $T$,$P$-violating matrix element presented in Eq. (25), uncertainties in the input data, the use of a hydrodynamic model and the use of static relations in the case of octupole vibrations, such as that in Eq. (7), lead to an overall uncertainty in our approach that may exceed a factor of 10.
@@
Calculations were not be able to be performed for 163Dy, 223Rn, 229Th and 233,235U (which are also theorised to have an octupole deformed nucleus or a soft octupole vibration mode) due to inconclusive $E1$ transition data in Ref. [48]. For a similar reason, we take the value of $d_{\text{int}}$ for 223Fr and 229Pa from other experiments, see Table 1.
@@
87 221Fr 5/2 0.14 11 0.13 0.25 0.20
87 223Fr 111$\lvert d_{\text{int}}\rvert$ obtained from Refs. [55, 56, 57]. 3/2 0.038 3.0 0.028 0.019 0.015
@@ V Effects of Enhanced nuclear moments in Molecules
In this work, we have performed a series of calculations characterising the enhancement of nuclear moments provided by the so called octupole mechanism. Octupole deformed nuclei have a low-lying energy level near the ground state, with opposing parity and identical spin. As such, the $T,P$-odd nuclear moments which are produced through the mixing of these doublet states by $T$,$P$-violating nuclear forces should be enhanced. Some enhancement may also appear in nuclei which are theorised to have a soft octupole vibration mode. Therefore, we have estimated the intrinsic nuclear electric dipole moment in a range of nuclei which are theorised to exhibit static octupole deformation, using half-life data for $E1$ transitions. In nuclei with a soft octupole vibration mode our estimates may be used as upper bounds.
These values were then used to present independent estimates of the octupole deformation parameter of each nucleus, as well as the intrinsic Schiff moment. The mixing of the rotational doublet states in these nuclei polarises the orientation of the nuclear axis along the nuclear spin, resulting in the manifestation of these intrinsic moments in the laboratory frame. This allows us to estimate the enhanced laboratory frame values of the nuclear EDM and Schiff moment in these nuclei. Our calculations indicate an enhancement of approximately 2-3 orders of magnitude of this collective Schiff moment when compared to the valence nucleon contribution to the nuclear Schiff moment in spherical nuclei. The nuclear EDM may be enhanced up to one order of magnitude.
@@
Our calculation of the nuclear EDM and Schiff moment may motivate the undertaking of resource-intensive nuclear many-body calculations for nuclei other than 225Ra and 153Eu (as well as corresponding atomic, molecular and solid state calculations) and has implications on the search for ultralight dark matter which produces an oscillating neutron EDM [32], nuclear EDM and Schiff moment [33]. In order to yield enhanced results, experiments measuring the axion field (or axion-like particle field) induced Schiff moment are seeking to probe crystals containing nuclei which benefit from the octupole mechanism. In particular, our calculation of the Schiff moment in 153Eu may be relevant to experimental groups currently probing crystals containing this isotope such as Refs. [40, 66].
@@ Acknowledgements
We are grateful to Alex Sushkov, Oleg Sushkov and Chang Zhou for useful discussions. The work was supported by the Australian Research Council Grant No. DP230101058.